# Emergent SU(3) topological system in a trimer SSH model

### v1

@@ III.2 Sub-lattice winding numbers as bulk-topological invariants
To elaborate on the above discussion, we again consider the trimer SSH chain with $u_{1}=v_{1}=0,~{}u=v=1$. In this case, the bulk energy band gap closes at ±$\pm 1$, therefore we choose the reference energies $E_{\rm b}=\pm 1$. The sub-lattice winding number $\nu_{\pm 1}^{\rm A}$ take the following form
@@
where $W=w_{1}+we^{-ik}$ is nothing but the off-diagonal element of $\mathcal{H}(k)$. The sub-lattice winding number $\nu_{\pm 1}^{\rm B}$ vanishes and $\nu_{\pm 1}^{\rm C}=-\nu_{\pm 1}^{\rm A}$ due to $\sum_{\alpha}\nu_{\pm 1}^{\alpha}=0$. It is straightforward to verify from Fig. 5 and Fig. 6 that $\nu^{\rm A}=\nu_{+1}^{\rm A}+\nu_{-1}^{\rm A}$ correctly counts the number of boundary modes at each end of the system. For example, in the case of $u_{1}=v_{1}=w_{1}=0$, for $w>u=v$ and $\nu^{\rm A}=2$ determines one boundary mode with energy $E_{\rm b}=-1$ and one boundary mode with energy $E_{\rm b}=+1$ at each end of the system. In the case of $u_{1}=v_{1}$, for $w>u=v,~{}w<w_{1}$, $\nu^{\rm A}=1$ determines the one boundary mode with energy $E_{\rm b}=-1$ at each end of the system and for $w>u=v,~{}w>w_{1}$, $\nu^{\rm A}=2$ determines one boundary mode with energy $E_{\rm b}=-1$ and one boundary mode with energy $E_{\rm b}=+1$ at each end of the system.
@@
The main advantage of this topological invariant is its analytical form and simple understanding of the topological phase transition, similar to the dimer SSH model [19]. When $k$ varies across the first BZ, the endpoint of the vector ${\bf W}_{\text{b}}$ traces out a circle of radius $w$ centered at $(w_{1}-E_{\text{b}})$ in the $W_{{\text{b}},x}$-$W_{{\text{b}},y}$ plane, as shown in the lower panels of Figs. 2 and 3. The gap closes at ${\bf W}_{\text{b}}=0$, which corresponds to the gap closing points $k=0,\pm\pi$ in $k$ space. The sub-lattice winding number will be one when the trajectory of ${\bf W}_{\text{b}}$ encloses the origin, otherwise zero. It is clear from Fig. 2 (d), (e), and (f) that the winding of the complex function $W_{\text{b}}$ is zero, ill-defined, and two for the trivial, critical, and non-trivial phases. The topological phase transition takes place at $w=u=v$ where the contour of $W_{\rm b}$ touches the origin which corresponds to the bulk-gap closing points at $k=0,\pi$. Figure 2 displays that tuning the parameter $w$ while fixing $u=v=1$ and $(u_{1},v_{1},w_{1})=(0,0,0)$ induces a topological phase transition characterized by the sub-lattice winding number from $\nu^{A}=0\leftrightarrow 2$.
@@
Similarly, it is clear from Fig. 3 (f)-(j) that the winding of $W_{\text{b}}$ is zero, ill-defined, one, one, and two, respectively. The topological phase transitions take place at $w=u=v,w<w_{1}$ where contour of $W+1$ touches the origin which corresponds to the bulk-gap closing at $k=\pm\pi$ and also at $w=u=v,w=w_{1}$ where the contour of $W-1$ touches the origin which corresponds to the bulk-gap closing at $k=\pm\pi$. Figure 3 displays a series of topological phase transitions characterized by the sub-lattice winding number from $\nu^{A}=0\leftrightarrow 1\leftrightarrow 2$, while tuning the parameter $w$ for fixed $u=v=1$ and $(u_{1},v_{1},w_{1})=(0,0,2)$.
IV Summary and Conclusion
@@
In this work, it is shown that the Bloch Hamiltonian of the trimer SSH model with the on-site potential energy and up to NNN hopping terms can be expressed in terms of all eight generators of the SU(3) group. The exact analytical expressions of the three dispersive energy bands and the associated eigenstates for arbitrary choices of the system parameters are provided. We discussed inversion symmetry, time-reversal symmetry, and some special chiral symmetry of the Hamiltonian. The system lacks full chiral symmetry, however, we find three different kinds of topological phase transitions while tuning the NN and NNN hopping terms. The topological phase transitions are characterized by the two bulk topological invariants, namely the Zak phase and the sub-lattice winding number. We established the bulk-boundary condition by computing the Zak phases ($\phi_{\lambda})$ for all the bands along with the boundary modes in the open boundary condition. The quantized change in two out of three Zak phases $\phi_{1}+\phi_{3}$ changing from $\phi_{1}+\phi_{3}=0\leftrightarrow 2\pi$ as well as $\phi_{1}+\phi_{3}=0\leftrightarrow\pi\leftrightarrow 2\pi$ characterize these topological phase transitions. We proposed a new topological invariant, the sub-lattice winding number which also characterizes the topological phase transitions. The exact analytical expressions of the sub-lattice winding number revealed that $\nu^{\alpha}$ changes from $\nu^{\alpha}=0\leftrightarrow 2$ as well as $\nu^{\alpha}=0\leftrightarrow 1\leftrightarrow 2$ across the topological phase transitions. The simple analytical method of calculating the sub-lattice winding number may help in understanding the topological phases of systems without chiral symmetry.
@@ ACKNOWLEDGEMENT
Appendix A Gell-Mann matrices and their basic properties
@@
The particles in the chain are spinless, and the time-reversal operator in the sublattice space is simply the complex conjugation operator $\mathcal{C}$. The three complex Gell-Mann matrices are odd under time reversal: $\mathcal{C}^{-1}\Lambda_{2,5,7}\mathcal{C}=-\Lambda_{2,5,7}$ and the remaining matrices are even under time reversal.
@@ Appendix C Different topological phases for $u_{1}=v_{1}\neq 0$ case
Figure C.7: Bulk-topological invariant: Here (a), (b), (c), (d), (e), and (f) describe the sum of the Zak phases $\phi_{1}+\phi_{3}$ (in units of $\pi$) in the $(w,w_{1})$ plane for $(u_{1},v_{1}):(0.2,0.2),(0.5,0.5),(1.0,1.0),(1.2,1.2),(1.5,1.5)$, and $(2.0,2.0)$, respectively. In all the plots here we have taken $u=v=1$
Figure C.7: Bulk-topological invariant: Here (a), (b), (c), (d), (e), and (f) describe the sum of the Zak phases $\phi_{1}+\phi_{3}$ (in units of $\pi$) in the $(w,w_{1})$ plane for $(u_{1},v_{1}):(0.2,0.2),(0.5,0.5),(1.0,1.0),(1.2,1.2),(1.5,1.5)$, and $(2.0,2.0)$, respectively. In all the plots here we have taken $u=v=1$
@@ Appendix D Sub-lattice winding number for the dimer SSH model
where $k$ denotes the Bloch wave vector and $d_{k}=t_{1}+t_{2}e^{-ik}$, with $t_{1}$ and $t_{2}$ represent nearest-neighbour intra-cell and inter-cell hopping parameters, respectively. This Hamiltonian respects time-reversal, particle-hole, chiral, and inversion symmetries. It supports edge state with energy $E_{\text{b}}=0$ at each end of the boundary.

### v2

@@ IV.1 Bulk case
To elaborate on the above discussion, we again consider the trimer SSH chain with $u_{1}=v_{1}=0,~{}u=v=1$. In this case, the bulk energy band gap closes at ±$\pm 1$ therefore we choose the reference energies $E_{\rm b}=\pm 1$. The sub-lattice winding number $\nu_{\pm 1}^{\rm A}$ take the following form
@@
where $W=w_{1}+we^{-ik}$ is nothing but the off-diagonal element of $\mathcal{H}(k)$. The sub-lattice winding number $\nu_{\pm 1}^{\rm B}$ vanishes and $\nu_{\pm 1}^{\rm C}=-\nu_{\pm 1}^{\rm A}$ due to $\sum_{\alpha}\nu_{\pm 1}^{\alpha}=0$. It is straightforward to verify from Fig. 6 and Fig. 7 that $\nu^{\rm A}=\nu_{+1}^{\rm A}+\nu_{-1}^{\rm A}$ correctly counts the number of boundary modes at each end of the system. For example, in the case of $u_{1}=v_{1}=w_{1}=0$, for $w>u=v$ and $\nu^{\rm A}=2$ determines one boundary mode with energy $E_{\rm b}=-1$ and one boundary mode with energy $E_{\rm b}=+1$ at each end of the system.
@@
One of the advantages of this topological invariant is its analytical form and simple understanding of the topological phase transition, similar to the dimer SSH model [23]. When $k$ varies across the first BZ, the endpoint of the vector ${\bf W}_{\text{b}}=W-E_{\text{b}}\equiv\hat{i}W_{{\text{b}},x}+\hat{j}W_{{\text{b}},y}$, with $W=w_{1}+we^{-ik},~{}W_{{\text{b}},x}={\rm Re}[W]-E_{\rm b},~{}W_{{\rm b},y}={\rm Im}[W]$, traces out a circle of radius $w$ centered at $(w_{1}-E_{\text{b}})$ in the $W_{{\text{b}},x}$-$W_{{\text{b}},y}$ plane, as shown in the lower panel of Fig. 2. The gap closes at ${\bf W}_{\text{b}}=0$, which corresponds to the gap closing points $k=0,\pm\pi$ in $k$ space. The sub-lattice winding number will be one when the trajectory of ${\bf W}_{\text{b}}$ encloses the origin, otherwise zero. It is clear from Fig. 2 (d), (e), and (f) that the winding of the complex function $W_{\text{b}}$ is zero, ill-defined, and two for the trivial, critical, and non-trivial phases. The topological phase transition takes place at $w=u=v$ where the contour of $W_{\rm b}$ touches the origin which corresponds to the bulk-gap closing points at $k=0,\pi$. Figure 2 displays that tuning the parameter $w$ while fixing $u=v=1$ and $(u_{1},v_{1},w_{1})=(0,0,0)$ induces a topological phase transition characterized by the sub-lattice winding number from $\nu^{A}=0\leftrightarrow 2$. Figure 5 (a) shows the variation of the sublattice winding number in the $(w,w_{1})$-plane, which shows that non-trivial values of the sublattice winding number coincide with the number of boundary modes in the open boundary condition.
Inversion asymmetric case: For the inversion asymmetric chain with $u\neq v,~{}w_{1}=u_{1}=v_{1}=0$, the boundary modes do not appear at the band-gap closing point; instead, they appear at the mirror-symmetric points at $w=\pm u$ and $w=\pm v$ with energies $E_{\rm b}=\pm u$ and $E_{\rm b}=\pm v$, respectively [48]. In this case, the bulk-boundary correspondence has been discussed in detail in Ref. [48] using the normalized sub-lattice Zak phase as a bulk topological invariant determined under the open boundary condition.
In this case, we also find that the sub-lattice winding number can still be quantized and successfully establish the bulk-boundary correspondence. Figure 5 (b) shows the variation of the sublattice winding number in the $(w,w_{1})$-plane. The non-trivial values of the sublattice winding number coincide with the number of boundary modes in the open boundary condition. To elaborate on the above findings, we show the energy band structure of a finite chain vs $w$, the sublattice winding number vs $w$, and the real space profile of the edge states in Fig. 8 (a), (b), and [(c)-(d)], respectively. We note here that the sublattice winding number can still be quantized and used to establish the bulk-boundary correspondence for $\Delta\neq 0$ with a more complicated phase diagram, which we do not discuss here.
IV.2 Finite chain
For $w>u=v$ and $w_{1}=0$, the Zak phases $\phi_{1},\phi_{3}$ are quantized to $\pi$ while $\phi_{2}$ vanishes. In the open boundary condition, there exists at each energy $E_{\text{b}}=\pm 1$, a single boundary mode localized at each end of the system [see Fig. 6 (a) ]. These results establish the bulk-boundary correspondence. In this case, the localization lengths for the boundary modes at energies $E_{\text{b}}=\pm 1$ are the same and are determined by the bulk energy gaps $\Delta_{21}$ and $\Delta_{32}$.
Figure 6 (a) and (b) show the variation of the band structure in the open boundary condition and $\phi_{1}+\phi_{3}$ with respect to $w$, respectively. Figures  6 (c) and (d) show the real space profile of the wave function amplitudes for a finite open chain corresponding to the boundary modes with energies $E_{\text{b}}=+1$ and $E_{\text{b}}=-1$, respectively.
In Ref. [48], it was shown that the following real space profile of edge states holds due to point chiral symmetry. In the inversion symmetric case, the wave function amplitudes vanish for all edge modes at one of three sub-lattice sites (at the A sub-lattice on the left and C sub-lattice on the right boundary). For the edge states with energies $E_{\rm b}=-u$ ($E_{\rm b}=u$), wave function amplitudes at the A and B sub-lattice sites are equal (equal and opposite). For the edge states with energies $E_{\rm b}=-v$ ($E_{\rm b}=v$), wave function amplitudes at the B and C sub-lattice sites are equal (equal and opposite). In the inversion asymmetric case, the boundary modes lose their inversion partner modes Figs. 8 (c) and (d). The boundary modes with energies $E_{\rm b}=\pm u$ and $E_{\rm b}=\pm v$ are localized on the left and right boundaries of the system, respectively. However amplitudes of the wave function for $E_{\rm b}=+u$ and $E_{\rm b}=-u$ (similarly for $E_{\rm b}=+v$ and $E_{\rm b}=-v$) do not coincide.
V Topological phases and bulk-boundary correspondence in the presence of NNN hopping
In this section, we discuss the bulk-boundary correspondence in the presence of NNN hopping terms ($u_{1},~{}v_{1},~{}w_{1}$) (Fig. 1).
V.1 Bulk
For finite NNN hopping ($w_{1}\neq 0$), the inversion symmetry enforces the hopping to satisfy the following criteria,
$\Delta=0,u=v,~{}\text{and}~{}u_{1}=v_{1}.$ (25)
The results for different sets of $u_{1}=v_{1}\neq 0$ is given in the Appendix C. In this case, the bulk band structure evolves according to Fig. 3 as we tune the inter-cell hopping $w$ for the fixed value of NNN hopping $w_{1}=2$. At $w=u=v$ and $w<w_{1}$, the band gap between the lower and middle bands closes at the $X$-point (See Fig.  3 (b)). Similarly, at $w>u=v$ and $w=w_{1}$, the band gap between the middle and the upper bands closes at the $X$-point (See Fig.  3 (d). We note here that in the presence of NNN hopping the only one of the band gaps $\Delta_{12}$ and $\Delta_{23}$ vanish either at $\Gamma$ or $X-$ points, which signals TPT-II and TPT-III, respectively (see Fig. C.10 in the Appendix.
We observe that the addition of Zak phases $\phi_{1}+\phi_{3}$ characterize the topological phase transitions and number of boundary modes at each end of the system. We note here that the quantized changes in the individual Zak phases $\phi_{1}$ and $\phi_{3}$ characterize the topological phase transitions (TPT)-(I and II) and (I and III), respectively. Instead, we use $\phi_{1}+\phi_{3}$ to convey the same information. A similar analysis holds for $u_{1}=v_{1}\neq 0$. However, the phase diagram looks more complex, which is given in Appendix C (See Fig. C.11).
In this case, we also find that the non-trivial values of the sublattice winding number coincide with the number of boundary modes in the open boundary condition, successfully establishing the BBC in the presence of NNN hopping. In the case of $u_{1}=v_{1}$, for $w>u=v,~{}w<w_{1}$, $\nu^{\rm A}=1$ [Eq. 23] determines the one boundary mode with energy $E_{\rm b}=-1$ at each end of the system and for $w>u=v,~{}w>w_{1}$, $\nu^{\rm A}=2$ determines one boundary mode with energy $E_{\rm b}=-1$ and one boundary mode with energy $E_{\rm b}=+1$ at each end of the system.
@@ V.1 Bulk
It is clear from Fig. 3 (f)-(j) that the winding of $W_{\text{b}}$ is zero, ill-defined, one, one, and two, respectively. The topological phase transitions take place at $w=u=v,w<w_{1}$ where contour of $W+1$ touches the origin, which corresponds to the bulk-gap closing at $k=\pm\pi$ and also at $w=u=v,w=w_{1}$ where the contour of $W-1$ touches the origin which corresponds to the bulk-gap closing at $k=\pm\pi$. Figure 3 displays a series of topological phase transitions characterized by the sub-lattice winding number from $\nu^{A}=0\leftrightarrow 1\leftrightarrow 2$, while tuning the parameter $w$ for fixed $u=v=1$ and $(u_{1},v_{1},w_{1})=(0,0,2)$. Fig. 5 (a) shows the variation of the sublattice winding number in the $(w,w_{1})$-plane. To elaborate on the above findings, we show the energy band structure of a finite chain vs $w$, the sublattice winding number vs $w$, and the real space profile of the edge states in Fig. 7 (a), (b), and [(c)-(d)].
Inversion asymmetric case: Similar to the case of no NNN hopping, in the inversion asymmetric case ($u\neq v$), we find that the sub-lattice winding number can still be quantized and successfully establishes the BBC in the presence of NNN hopping. For $w_{1}\neq 0$, the boundary modes appear at $w=\pm(w_{1}\pm u)$ and $w=\pm(w_{1}\pm v)$; however, their energies $E_{\rm b}=\pm u$ and $E_{\rm b}=\pm v$ (which remain the same), respectively. Fig. 5 (b) shows the variation of the sublattice winding number in the $(w,w_{1})$-plane. The non-trivial values of the sublattice winding number coincide with the number of boundary modes in the open boundary condition. To elaborate on the above findings, we show the energy band structure of a finite chain vs $w$, the sublattice winding number vs $w$, and the real space profile of the edge states in Fig. 9 (a), (b), and [(c)-(d)].
V.2 Finite chain
Figs. 7 (a) and (b) show the variation of the band structure in the open boundary condition and $\phi_{1}+\phi_{3}$ with respect to $w$, respectively. Figs. 7 (c) and (d) show the real space profile of wave function amplitudes for a finite open chain corresponding to the boundary modes with energies $E_{\text{b}}=+1$ and $E_{\text{b}}=-1$, respectively. Accordingly, in the open boundary condition, there exist at each energy $E_{\text{b}}=\pm 1$, a single boundary mode localized at each end of the system. In this case the localization lengths for the boundary modes at energies $E_{\text{b}}=\pm 1$ are different and are determined by the bulk energy gap $\Delta_{21}$ ($\Delta_{32}$) for negative (positive) energy boundary modes.
In the inversion symmetric case [Figs. 7 (c) and (d)], for all edge modes, the wave function amplitudes vanish at one of three sub-lattice sites (at A sub-lattice on the left and C sub-lattice on the right boundary). For the edge states with energies $E_{\rm b}=-u$ ($E_{\rm b}=u$), wave function amplitudes at the A and B sub-lattice sites are equal (equal and opposite). For the edge states with energies $E_{\rm b}=-v$ ($E_{\rm b}=v$), wave function amplitudes at the B and C sub-lattice sites are equal (equal and opposite).
Inversion asymmetric case: In this case, the boundary modes lose their inversion partner modes Figs. 9 (c) and (d). The boundary modes with energies $E_{\rm b}=\pm u$ and $E_{\rm b}=\pm v$ are localized on the left and right boundaries of the system, respectively. However amplitudes of the wave function for $E_{\rm b}=+u$ and $E_{\rm b}=-u$ (similarly for $E_{\rm b}=+v$ and $E_{\rm b}=-v$) do not coincide. In Ref. [48], it was also argued that these real space profile of edge states holds due to point chiral symmetry, which is broken for $w_{1}\neq 0$). It is unclear which symmetry enforces the properties of the edge states and quantization of the sublattice winding number even in the absence of inversion symmetry ($u\neq v$) and point chiral symmetry ($w_{1}\neq 0$). To address these queries further, the studies [57, 58] of the bulk-boundary correspondences in systems with full or partial symmetry-breaking perturbations using non-Bloch band theory may be useful.
VI Summary and Conclusion
@@
In this work, it is shown that the Bloch Hamiltonian of the trimer SSH model with the on-site potential energy and up to NNN hopping terms can be expressed in terms of all eight generators of the SU(3) group. The exact analytical expressions of the three dispersive energy bands and the associated eigenstates for arbitrary choices of the system parameters are provided. We discussed inversion symmetry, time-reversal symmetry, and some special chiral symmetry of the Hamiltonian. The system lacks full chiral symmetry, however, we find three different kinds of topological phase transitions while tuning the NN and NNN hopping terms. The topological phase transitions are characterized by the two bulk topological invariants, namely the Zak phase and the sub-lattice winding number. We established the bulk-boundary correspondence by computing the Zak phases ($\phi_{\lambda})$ for all the bands along with the boundary modes in the open boundary condition. The quantized change in two out of three Zak phases $\phi_{1}+\phi_{3}$ changing from $\phi_{1}+\phi_{3}=0\leftrightarrow 2\pi$ as well as $\phi_{1}+\phi_{3}=0\leftrightarrow\pi\leftrightarrow 2\pi$ characterize these topological phase transitions. We proposed a new topological invariant, the sub-lattice winding number which also characterizes the topological phase transitions. The exact analytical expressions of the sub-lattice winding number revealed that $\nu^{\alpha}$ changes from $\nu^{\alpha}=0\leftrightarrow 2$ as well as $\nu^{\alpha}=0\leftrightarrow 1\leftrightarrow 2$ across the topological phase transitions. The simple analytical method of calculating the sub-lattice winding number and its success in establishing the bulk-boundary correspondence in the inversion asymmetric system may help understand the bulk-boundary correspondence in systems without chiral and inversion symmetry.
@@ ACKNOWLEDGEMENT
Appendix A Gell-Mann matrices and a couple of basic properties
@@
The particles in the chain are spinless, and the time-reversal operator in the sub-lattice space is simply the complex conjugation operator $\mathcal{C}$. The three complex Gell-Mann matrices are odd under time reversal: $\mathcal{C}^{-1}\Lambda_{2,5,7}\mathcal{C}=-\Lambda_{2,5,7}$ and the remaining matrices are even under time reversal.
@@ Appendix C Different topological phases for $u_{1}=v_{1}\neq 0$ case
Figure C.10: Bloch band structure: In all the plots here, we have taken $u=v=1,~{}u_{1}=v_{1}=0$.
Figure C.11: Bulk-topological invariant: Here (a), (b), (c), (d), (e), and (f) describe the sum of the Zak phases $\phi_{1}+\phi_{3}$ (in units of $\pi$) in the $(w,w_{1})$ plane for $(u_{1},v_{1}):(0.2,0.2),(0.5,0.5),(1.0,1.0),(1.2,1.2),(1.5,1.5)$, and $(2.0,2.0)$, respectively. In all the plots here we have taken $u=v=1$.
@@ Appendix D Sub-lattice winding number for the dimer SSH model
where $k$ denotes the Bloch wave vector and $d_{k}=t_{1}+t_{2}e^{-ik}$, with $t_{1}$ and $t_{2}$ represent nearest-neighbour intra-cell and inter-cell hopping parameters, respectively. This Hamiltonian respects time-reversal, inversion, chiral, and particle-hole symmetries. It supports boundary mode with energy $E_{\text{b}}=0$ at each end of the system when $t_{2}>t_{1}$.
@@
Appendix E Real space profile of edge states and their connections to topological invariants
Consider the eigenvalue equation in the open boundary condition,
$H|\psi\rangle=\pm u|\psi\rangle,$ (63)
with
$|\psi\rangle=(\psi_{1\rm A},\psi_{1\rm B},\psi_{1\rm C},\psi_{2\rm A},\psi_{2\rm B},\psi_{2\rm C},...,\psi_{N\rm A},\psi_{N\rm B},\psi_{N\rm C})^{T}$ and
$H=\begin{pmatrix}0&u&w_{1}&0&0&0&0&0&0\\ u&0&v&u_{1}&0&0&0&0&0\\ w_{1}&v&0&w&v_{1}&0&0&0&0\\ 0&u_{1}&w&0&u&w_{1}&0&0&0\\ 0&0&v_{1}&u&0&v&u_{1}&0&0\\ 0&0&0&w_{1}&v&0&w&v_{1}&0\\ 0&0&0&0&u_{1}&w&0&u&w_{1}\\ 0&0&0&0&0&v_{1}&u&0&v\\ 0&0&0&0&0&0&w_{1}&v&0\end{pmatrix}$ (64)
which immediately implies the following set of equations
$\mp u\psi_{1\rm A}+u\psi_{1\rm B}+w_{1}\psi_{1\rm C}$ $=0,$ (65)
$u\psi_{1\rm A}\mp u\psi_{1\rm B}+v\psi_{1\rm C}+u_{1}\psi_{2\rm A}$ $=0,...$ (66)
For $u_{1}=0$, solving the above two equations immediately gives rise to
(i) $\psi_{1\rm C}=0,~{}\psi_{1\rm A}=\psi_{1\rm B}$ for $E_{\rm b}=-u$,
(ii) $\psi_{1\rm C}=0,~{}\psi_{1\rm A}=-\psi_{1\rm B}$ for $E_{\rm b}=u$.
Similarly, the eigenvalue equation for $E_{\rm b}=\pm v$ leads to
$w_{1}\psi_{N\rm A}+v\psi_{N\rm B}\mp v\psi_{N\rm C}$ $=0,$ (67)
$v_{1}\psi_{(N-1)\rm C}+u\psi_{N\rm A}\mp v\psi_{N\rm B}+v\psi_{N\rm C}$ $=0,...$ (68)
For $v_{1}=0$, solving the above two equations immediately gives rise to
(i) $\psi_{N\rm A}=0,~{}\psi_{N\rm B}=\psi_{N\rm C}$ for $E_{\rm b}=-v$,
(ii) $\psi_{N\rm A}=0,~{}\psi_{N\rm B}=-\psi_{N\rm C}$ for $E_{\rm b}=v$.
Therefore, we conclude that the wave function for edge states with energies $E_{\rm b}=-u(u)$ and $E_{\rm b}=-v(v)$ have equal (equal and opposite) support on even and odd sub-lattice sites even in the presence of next-nearest neighbor hopping $w_{1}\neq 0$ where the point chiral and or inversion symmetries are lost.